OpenStudy (anonymous):

Simplify the expression. (-6i) (-6i)

Parth (parthkohli):

\( \color{Black}{\Rightarrow -6 \times -6 \times i^2 }\)

OpenStudy (lgbasallote):

-6i x -6i = 6i^2 i^2 = -1 therefore 6i^2 = -6

OpenStudy (anonymous):

\[36i^2\] \[(36)*(-1)\] =36

Parth (parthkohli):

What is \(i^2\)?

OpenStudy (lgbasallote):

imean 36 lol

OpenStudy (anonymous):

-36

OpenStudy (lgbasallote):

im getting old for this thing =)))

OpenStudy (anonymous):

i^2 is an imaginary number @ParthKohli

Parth (parthkohli):

I know... I asked him :/

OpenStudy (anonymous):

Oh sorry...

Parth (parthkohli):

And \(i\) is imaginary, not \(i^2\) :)

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

Well, yes and you are welcome tryder!